OpenStudy (anonymous):

5x^3y+xy^4+2xy ---------------- xy

OpenStudy (anonymous):

y^3+5x^2+2

OpenStudy (anonymous):

can i put it like this 5x^3+y^3+2

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well, no. The reason is because you are using 5x^3 and you need 5x^2 instead. :)

OpenStudy (anonymous):

x to th second rite

OpenStudy (anonymous):

You have to break up the sum in the numerator into different fractions: \[\frac{5x^3y+xy^4+2xy}{xy} = \frac{5x^3y}{xy}+\frac{xy^4}{xy}+\frac{2xy}{xy}\]

OpenStudy (anonymous):

Yes, x to the 2nd.

OpenStudy (anonymous):

Then you simplify from there the way you were before.

OpenStudy (anonymous):

thanx